# A robust mixed-effects quantile regression model using generalized Laplace mixtures to handle outliers and skewness

### v1

@@ Acknowledgements
AL GAL cGAL
Quantilea Parameter Est. 95% HPD Est. 95% HPD Est. 95% HPD
$0.05$ $\beta_{1}$ $10.888$ [10.473; $11.311$] $10.708$ [10.317; $11.120$] $10.766$ [10.375; $11.169$]
$\beta_{2}$ $0.319$ [0.280; $0.363$] $0.313$ [0.276; $0.355$] $0.318$ [0.279; $0.361$]
$\beta_{3}$ $5.590$ [4.912; $6.217$] $5.382$ [4.740; $6.015$] $5.546$ [4.845; $6.224$]
$\beta_{4}$ $-0.003$ [-0.021; $0.013$] $-0.005$ [-0.020; $0.011$] $-0.002$ [-0.019; $0.013$]
$\beta_{5}$ $0.004$ [-0.003; $0.010$] $0.004$ [-0.001; $0.010$] $0.003$ [-0.003; $0.009$]
$\sigma$ $0.021$ [0.018; $0.024$] $0.067$ [0.055; $0.078$] $0.057$ [0.036; $0.071$]
$\gamma$ $2.847$ [1.968; $3.824$] $4.058$ [2.225; $6.975$]
$\alpha$ $0.018$ [0.002; $0.038$]
$\omega_{11}$ $0.929$ [0.452; $1.600$] $1.139$ [0.515; $2.210$] $1.094$ [0.484; $1.987$]
$\omega_{21}$ $2.042$ [-2.583; $8.240$] $3.393$ [-3.734; $15.857$] $2.958$ [-3.758; $13.770$]
$\omega_{22}$ $126.647$ [45.138; $273.301$] $182.982$ [41.391; $495.406$] $175.010$ [57.743; $442.609$]
$\omega_{31}$ $-0.374$ [-0.807; $-0.056$] $-0.469$ [-1.102; $-0.092$] $-0.465$ [-1.045; $-0.087$]
$\omega_{32}$ $-0.377$ [-5.359; $4.413$] $-1.594$ [-10.095; $5.187$] $-1.945$ [-10.519; $3.920$]
$\omega_{33}$ $0.647$ [0.329; $1.079$] $0.728$ [0.346; $1.262$] $0.719$ [0.333; $1.231$]
$\omega_{41}$ $14.600$ [-8.951; $44.343$] $15.825$ [-15.627; $51.615$] $13.007$ [-12.288; $42.097$]
$\omega_{42}$ $-106.832$ [-534.535; $273.684$] $-165.956$ [-823.793; $408.179$] $-176.357$ [-774.993; $198.428$]
$\omega_{43}$ $-24.531$ [-51.701; $-4.272$] $-25.121$ [-51.457; $-2.802$] $-22.975$ [-45.799; $-4.021$]
$\omega_{44}$ $3550.850$ [1563.010; $7617.490$] $3775.830$ [1668.130; $7478.400$] $3171.280$ [1607.770; $5613.480$]
$0.25$ $\beta_{1}$ $11.200$ [10.806; $11.592$] $11.204$ [10.813; $11.612$] $11.264$ [10.846; $11.642$]
$\beta_{2}$ $0.313$ [0.277; $0.354$] $0.312$ [0.273; $0.352$] $0.319$ [0.282; $0.359$]
$\beta_{3}$ $5.953$ [5.261; $6.573$] $5.904$ [5.313; $6.553$] $6.098$ [5.396; $6.806$]
$\beta_{4}$ $-0.004$ [-0.018; $0.011$] $-0.007$ [-0.022; $0.008$] $-0.002$ [-0.020; $0.014$]
$\beta_{5}$ $0.004$ [-0.001; $0.009$] $0.004$ [-0.001; $0.010$] $0.003$ [-0.002; $0.009$]
$\sigma$ $0.093$ [0.081; $0.107$] $0.111$ [0.095; $0.131$] $0.097$ [0.041; $0.121$]
$\gamma$ $0.473$ [0.063; $0.833$] $0.745$ [0.129; $2.113$]
$\alpha$ $0.022$ [0.002; $0.051$]
$\omega_{11}$ $1.112$ [0.506; $2.040$] $1.174$ [0.532; $2.177$] $1.115$ [0.543; $2.030$]
$\omega_{21}$ $2.896$ [-4.712; $16.280$] $3.385$ [-5.005; $19.242$] $2.688$ [-6.698; $16.022$]
$\omega_{22}$ $206.009$ [61.341; $610.099$] $218.843$ [57.340; $719.243$] $213.755$ [61.322; $782.779$]
$\omega_{31}$ $-0.447$ [-1.019; $-0.043$] $-0.487$ [-1.108; $-0.079$] $-0.465$ [-1.085; $-0.087$]
$\omega_{32}$ $-1.145$ [-11.137; $7.147$] $-1.869$ [-14.249; $5.966$] $-1.524$ [-12.826; $10.053$]
$\omega_{33}$ $0.729$ [0.339; $1.250$] $0.753$ [0.344; $1.353$] $0.755$ [0.330; $1.328$]
$\omega_{41}$ $17.577$ [-10.986; $50.696$] $18.502$ [-11.716; $52.047$] $12.994$ [-16.221; $40.612$]
$\omega_{42}$ $-181.731$ [-935.112; $286.107$] $-194.774$ [-1036.640; $330.180$] $-222.341$ [-1062.980; $360.416$]
$\omega_{43}$ $-27.220$ [-55.447; $-6.385$] $-26.756$ [-55.040; $-5.211$] $-24.384$ [-50.962; $-3.354$]
$\omega_{44}$ $3734.935$ [1592.020; $7133.150$] $3600.175$ [1598.020; $6855.200$] $3220.385$ [1497.850; $6166.040$]
$0.5$ $\beta_{1}$ $11.614$ [11.216; $12.011$] $11.572$ [11.216; $11.993$] $11.591$ [11.220; $11.980$]
$\beta_{2}$ $0.318$ [0.281; $0.359$] $0.314$ [0.275; $0.356$] $0.318$ [0.280; $0.359$]
$\beta_{3}$ $6.327$ [5.698; $6.913$] $6.289$ [5.684; $6.897$] $6.380$ [5.729; $7.047$]
$\beta_{4}$ $-0.010$ [-0.024; $0.006$] $-0.006$ [-0.023; $0.008$] $-0.005$ [-0.022; $0.012$]
$\beta_{5}$ $0.005$ [-0.001; $0.010$] $0.004$ [-0.001; $0.010$] $0.004$ [-0.001; $0.010$]
$\sigma$ $0.126$ [0.109; $0.144$] $0.112$ [0.020; $0.137$] $0.102$ [0.039; $0.129$]
$\gamma$ $-0.326$ [-0.983; $-0.057$] $-0.242$ [-0.836; $0.097$]
$\alpha$ $0.018$ [0.001; $0.045$]
$\omega_{11}$ $1.201$ [0.559; $2.336$] $1.180$ [0.504; $2.259$] $1.112$ [0.543; $2.066$]
$\omega_{21}$ $3.293$ [-5.487; $17.590$] $3.526$ [-4.655; $18.509$] $2.735$ [-6.302; $16.953$]
$\omega_{22}$ $206.829$ [64.149; $684.648$] $218.613$ [61.656; $657.225$] $214.907$ [58.781; $738.424$]
$\omega_{31}$ $-0.487$ [-1.140; $-0.071$] $-0.491$ [-1.147; $-0.058$] $-0.465$ [-1.055; $-0.069$]
$\omega_{32}$ $-1.710$ [-13.728; $7.355$] $-1.855$ [-12.912; $6.737$] $-1.912$ [-13.632; $9.802$]
$\omega_{33}$ $0.775$ [0.376; $1.435$] $0.768$ [0.357; $1.382$] $0.744$ [0.340; $1.314$]
$\omega_{41}$ $18.944$ [-10.219; $51.400$] $17.725$ [-13.623; $50.222$] $15.522$ [-10.741; $46.174$]
$\omega_{42}$ $-165.159$ [-949.283; $332.340$] $-202.437$ [-989.776; $346.832$] $-218.613$ [-1114.230; $273.108$]
$\omega_{43}$ $-26.966$ [-56.662; $-6.934$] $-27.142$ [-55.074; $-5.559$] $-24.462$ [-49.770; $-4.232$]
$\omega_{44}$ $3289.345$ [1451.350; $5738.650$] $3568.830$ [1648.540; $6723.000$] $3166.495$ [1543.930; $5895.370$]
$0.75$ $\beta_{1}$ $12.152$ [11.737; $12.546$] $12.071$ [11.691; $12.464$] $12.032$ [11.603; $12.402$]
$\beta_{2}$ $0.340$ [0.300; $0.388$] $0.318$ [0.278; $0.365$] $0.323$ [0.282; $0.368$]
$\beta_{3}$ $7.005$ [6.334; $7.574$] $6.766$ [6.116; $7.404$] $6.767$ [6.157; $7.384$]
$\beta_{4}$ $-0.016$ [-0.034; $0.001$] $-0.011$ [-0.027; $0.006$] $-0.007$ [-0.022; $0.009$]
$\beta_{5}$ $0.006$ [-0.000; $0.013$] $0.005$ [-0.001; $0.011$] $0.004$ [-0.002; $0.010$]
$\sigma$ $0.095$ [0.081; $0.109$] $0.104$ [0.072; $0.130$] $0.085$ [0.043; $0.113$]
$\gamma$ $-1.240$ [-1.804; $-0.741$] $-1.408$ [-2.100; $-0.854$]
$\alpha$ $0.015$ [0.001; $0.035$]
$\omega_{11}$ $1.170$ [0.503; $2.123$] $1.232$ [0.539; $2.447$] $1.136$ [0.544; $2.105$]
$\omega_{21}$ $1.698$ [-9.533; $24.553$] $3.447$ [-4.343; $17.750$] $2.972$ [-3.983; $14.787$]
$\omega_{22}$ $156.002$ [30.900; $2301.640$] $193.226$ [48.892; $570.078$] $186.693$ [59.594; $472.967$]
$\omega_{31}$ $-0.446$ [-1.040; $-0.029$] $-0.498$ [-1.261; $-0.066$] $-0.462$ [-1.064; $-0.075$]
$\omega_{32}$ $-0.271$ [-9.905; $24.276$] $-2.358$ [-12.522; $7.086$] $-2.335$ [-10.780; $4.892$]
$\omega_{33}$ $0.733$ [0.336; $1.416$] $0.770$ [0.380; $1.468$] $0.736$ [0.367; $1.339$]
$\omega_{41}$ $18.636$ [-3.281; $44.942$] $18.428$ [-8.155; $53.426$] $14.099$ [-12.459; $42.389$]
$\omega_{42}$ $-37.728$ [-730.643; $916.398$] $-159.418$ [-857.598; $280.474$] $-204.434$ [-837.495; $225.430$]
[…]

### v2

@@ Acknowledgements
[…]
Quantilea Par. Est. 95% HPD Est. 95% HPD Est. 95% HPD Est. 95% HPD
$0.05$ $\beta_{1}$ $10.888$ [10.473; $11.311$] $10.708$ [10.317; $11.120$] $10.766$ [10.375; $11.169$] $10.761$ [10.386; $11.189$]
$\beta_{2}$ $0.319$ [0.280; $0.363$] $0.313$ [0.276; $0.355$] $0.318$ [0.279; $0.361$] $0.319$ [0.282; $0.363$]
$\beta_{3}$ $5.590$ [4.912; $6.217$] $5.382$ [4.740; $6.015$] $5.546$ [4.845; $6.224$] $5.555$ [4.878; $6.216$]
$\beta_{4}$ $-0.003$ [-0.021; $0.013$] $-0.005$ [-0.020; $0.011$] $-0.002$ [-0.019; $0.013$] $-0.002$ [-0.018; $0.014$]
$\beta_{5}$ $0.004$ [-0.003; $0.010$] $0.004$ [-0.001; $0.010$] $0.003$ [-0.003; $0.009$] $0.003$ [-0.002; $0.009$]
$\sigma$ $0.021$ [0.018; $0.024$] $0.067$ [0.055; $0.078$] $0.057$ [0.036; $0.071$] $0.058$ [0.038; $0.070$]
$\gamma$ $2.847$ [1.968; $3.824$] $4.058$ [2.225; $6.975$] $3.938$ [2.190; $6.720$]
$\alpha$ $0.018$ [0.002; $0.038$] $0.015$ [0.002; $0.037$]
$\omega_{11}$ $0.929$ [0.452; $1.600$] $1.139$ [0.515; $2.210$] $1.094$ [0.484; $1.987$] $1.097$ [0.451; $1.998$]
$\omega_{21}$ $2.042$ [-2.583; $8.240$] $3.393$ [-3.734; $15.857$] $2.958$ [-3.758; $13.770$] $2.866$ [-4.008; $13.806$]
$\omega_{22}$ $126.647$ [45.138; $273.301$] $182.982$ [41.391; $495.406$] $175.010$ [57.743; $442.609$] $175.704$ [60.605; $448.345$]
$\omega_{31}$ $-0.374$ [-0.807; $-0.056$] $-0.469$ [-1.102; $-0.092$] $-0.465$ [-1.045; $-0.087$] $-0.463$ [-1.040; $-0.094$]
$\omega_{32}$ $-0.377$ [-5.359; $4.413$] $-1.594$ [-10.095; $5.187$] $-1.945$ [-10.519; $3.920$] $-2.053$ [-10.437; $4.425$]
$\omega_{33}$ $0.647$ [0.329; $1.079$] $0.728$ [0.346; $1.262$] $0.719$ [0.333; $1.231$] $0.713$ [0.356; $1.217$]
$\omega_{41}$ $14.600$ [-8.951; $44.343$] $15.825$ [-15.627; $51.615$] $13.007$ [-12.288; $42.097$] $13.095$ [-12.239; $40.346$]
$\omega_{42}$ $-106.832$ [-534.535; $273.684$] $-165.956$ [-823.793; $408.179$] $-176.357$ [-774.993; $198.428$] $-186.963$ [-742.759; $216.296$]
$\omega_{43}$ $-24.531$ [-51.701; $-4.272$] $-25.121$ [-51.457; $-2.802$] $-22.975$ [-45.799; $-4.021$] $-22.997$ [-45.336; $-3.144$]
$\omega_{44}$ $3550.850$ [1563.010; $7617.490$] $3775.830$ [1668.130; $7478.400$] $3171.280$ [1607.770; $5613.480$] $3140.775$ [1540.320; $5472.120$]
$0.25$ $\beta_{1}$ $11.200$ [10.806; $11.592$] $11.204$ [10.813; $11.612$] $11.264$ [10.846; $11.642$] $11.243$ [10.843; $11.646$]
$\beta_{2}$ $0.313$ [0.277; $0.354$] $0.312$ [0.273; $0.352$] $0.319$ [0.282; $0.359$] $0.319$ [0.282; $0.360$]
$\beta_{3}$ $5.953$ [5.261; $6.573$] $5.904$ [5.313; $6.553$] $6.098$ [5.396; $6.806$] $6.071$ [5.440; $6.758$]
$\beta_{4}$ $-0.004$ [-0.018; $0.011$] $-0.007$ [-0.022; $0.008$] $-0.002$ [-0.020; $0.014$] $-0.002$ [-0.018; $0.013$]
$\beta_{5}$ $0.004$ [-0.001; $0.009$] $0.004$ [-0.001; $0.010$] $0.003$ [-0.002; $0.009$] $0.003$ [-0.002; $0.010$]
$\sigma$ $0.093$ [0.081; $0.107$] $0.111$ [0.095; $0.131$] $0.097$ [0.041; $0.121$] $0.098$ [0.052; $0.123$]
$\gamma$ $0.473$ [0.063; $0.833$] $0.745$ [0.129; $2.113$] $0.698$ [0.053; $1.876$]
$\alpha$ $0.022$ [0.002; $0.051$] $0.018$ [0.000; $0.044$]
$\omega_{11}$ $1.112$ [0.506; $2.040$] $1.174$ [0.532; $2.177$] $1.115$ [0.543; $2.030$] $1.109$ [0.538; $2.050$]
$\omega_{21}$ $2.896$ [-4.712; $16.280$] $3.385$ [-5.005; $19.242$] $2.688$ [-6.698; $16.022$] $2.712$ [-5.798; $15.642$]
$\omega_{22}$ $206.009$ [61.341; $610.099$] $218.843$ [57.340; $719.243$] $213.755$ [61.322; $782.779$] $208.914$ [51.528; $666.532$]
$\omega_{31}$ $-0.447$ [-1.019; $-0.043$] $-0.487$ [-1.108; $-0.079$] $-0.465$ [-1.085; $-0.087$] $-0.453$ [-1.056; $-0.076$]
$\omega_{32}$ $-1.145$ [-11.137; $7.147$] $-1.869$ [-14.249; $5.966$] $-1.524$ [-12.826; $10.053$] $-1.748$ [-12.308; $7.703$]
$\omega_{33}$ $0.729$ [0.339; $1.250$] $0.753$ [0.344; $1.353$] $0.755$ [0.330; $1.328$] $0.745$ [0.342; $1.304$]
$\omega_{41}$ $17.577$ [-10.986; $50.696$] $18.502$ [-11.716; $52.047$] $12.994$ [-16.221; $40.612$] $13.213$ [-12.419; $42.225$]
$\omega_{42}$ $-181.731$ [-935.112; $286.107$] $-194.774$ [-1036.640; $330.180$] $-222.341$ [-1062.980; $360.416$] $-230.349$ [-1126.830; $318.367$]
$\omega_{43}$ $-27.220$ [-55.447; $-6.385$] $-26.756$ [-55.040; $-5.211$] $-24.384$ [-50.962; $-3.354$] $-23.757$ [-47.201; $-1.912$]
$\omega_{44}$ $3734.935$ [1592.020; $7133.150$] $3600.175$ [1598.020; $6855.200$] $3220.385$ [1497.850; $6166.040$] $3230.330$ [1495.320; $5815.670$]
$0.5$ $\beta_{1}$ $11.614$ [11.216; $12.011$] $11.572$ [11.216; $11.993$] $11.591$ [11.220; $11.980$] $11.595$ [11.187; $11.977$]
$\beta_{2}$ $0.318$ [0.281; $0.359$] $0.314$ [0.275; $0.356$] $0.318$ [0.280; $0.359$] $0.317$ [0.278; $0.355$]
$\beta_{3}$ $6.327$ [5.698; $6.913$] $6.289$ [5.684; $6.897$] $6.380$ [5.729; $7.047$] $6.366$ [5.723; $7.056$]
$\beta_{4}$ $-0.010$ [-0.024; $0.006$] $-0.006$ [-0.023; $0.008$] $-0.005$ [-0.022; $0.012$] $-0.006$ [-0.024; $0.009$]
$\beta_{5}$ $0.005$ [-0.001; $0.010$] $0.004$ [-0.001; $0.010$] $0.004$ [-0.001; $0.010$] $0.004$ [-0.001; $0.010$]
$\sigma$ $0.126$ [0.109; $0.144$] $0.112$ [0.020; $0.137$] $0.102$ [0.039; $0.129$] $0.102$ [0.022; $0.127$]
$\gamma$ $-0.326$ [-0.983; $-0.057$] $-0.242$ [-0.836; $0.097$] $-0.260$ [-0.961; $0.045$]
$\alpha$ $0.018$ [0.001; $0.045$] $0.015$ [0.000; $0.041$]
$\omega_{11}$ $1.201$ [0.559; $2.336$] $1.180$ [0.504; $2.259$] $1.112$ [0.543; $2.066$] $1.131$ [0.527; $2.006$]
$\omega_{21}$ $3.293$ [-5.487; $17.590$] $3.526$ [-4.655; $18.509$] $2.735$ [-6.302; $16.953$] $2.668$ [-5.229; $15.098$]
$\omega_{22}$ $206.829$ [64.149; $684.648$] $218.613$ [61.656; $657.225$] $214.907$ [58.781; $738.424$] $206.373$ [59.435; $591.464$]
$\omega_{31}$ $-0.487$ [-1.140; $-0.071$] $-0.491$ [-1.147; $-0.058$] $-0.465$ [-1.055; $-0.069$] $-0.459$ [-1.018; $-0.081$]
$\omega_{32}$ $-1.710$ [-13.728; $7.355$] $-1.855$ [-12.912; $6.737$] $-1.912$ [-13.632; $9.802$] $-1.687$ [-11.740; $6.298$]
$\omega_{33}$ $0.775$ [0.376; $1.435$] $0.768$ [0.357; $1.382$] $0.744$ [0.340; $1.314$] $0.736$ [0.362; $1.302$]
$\omega_{41}$ $18.944$ [-10.219; $51.400$] $17.725$ [-13.623; $50.222$] $15.522$ [-10.741; $46.174$] $14.900$ [-11.623; $42.636$]
$\omega_{42}$ $-165.159$ [-949.283; $332.340$] $-202.437$ [-989.776; $346.832$] $-218.613$ [-1114.230; $273.108$] $-201.720$ [-850.471; $286.951$]
$\omega_{43}$ $-26.966$ [-56.662; $-6.934$] $-27.142$ [-55.074; $-5.559$] $-24.462$ [-49.770; $-4.232$] $-24.282$ [-48.854; $-5.650$]
$\omega_{44}$ $3289.345$ [1451.350; $5738.650$] $3568.830$ [1648.540; $6723.000$] $3166.495$ [1543.930; $5895.370$] $3174.795$ [1464.620; $5535.030$]
$0.75$ $\beta_{1}$ $12.152$ [11.737; $12.546$] $12.071$ [11.691; $12.464$] $12.032$ [11.603; $12.402$] $12.036$ [11.674; $12.438$]
$\beta_{2}$ $0.340$ [0.300; $0.388$] $0.318$ [0.278; $0.365$] $0.323$ [0.282; $0.368$] $0.326$ [0.284; $0.370$]
$\beta_{3}$ $7.005$ [6.334; $7.574$] $6.766$ [6.116; $7.404$] $6.767$ [6.157; $7.384$] $6.785$ [6.155; $7.451$]
$\beta_{4}$ $-0.016$ [-0.034; $0.001$] $-0.011$ [-0.027; $0.006$] $-0.007$ [-0.022; $0.009$] $-0.007$ [-0.023; $0.011$]
$\beta_{5}$ $0.006$ [-0.000; $0.013$] $0.005$ [-0.001; $0.011$] $0.004$ [-0.002; $0.010$] $0.004$ [-0.001; $0.011$]
$\sigma$ $0.095$ [0.081; $0.109$] $0.104$ [0.072; $0.130$] $0.085$ [0.043; $0.113$] $0.085$ [0.043; $0.112$]
$\gamma$ $-1.240$ [-1.804; $-0.741$] $-1.408$ [-2.100; $-0.854$] $-1.399$ [-2.066; $-0.788$]
$\alpha$ $0.015$ [0.001; $0.035$] $0.013$ [0.001; $0.033$]
$\omega_{11}$ $1.170$ [0.503; $2.123$] $1.232$ [0.539; $2.447$] $1.136$ [0.544; $2.105$] $1.147$ [0.502; $2.130$]
$\omega_{21}$ $1.698$ [-9.533; $24.553$] $3.447$ [-4.343; $17.750$] $2.972$ [-3.983; $14.787$] $2.912$ [-3.855; $14.449$]
$\omega_{22}$ $156.002$ [30.900; $2301.640$] $193.226$ [48.892; $570.078$] $186.693$ [59.594; $472.967$] $180.619$ [57.655; $461.854$]
$\omega_{31}$ $-0.446$ [-1.040; $-0.029$] $-0.498$ [-1.261; $-0.066$] $-0.462$ [-1.064; $-0.075$] $-0.473$ [-1.102; $-0.076$]
$\omega_{32}$ $-0.271$ [-9.905; $24.276$] $-2.358$ [-12.522; $7.086$] $-2.335$ [-10.780; $4.892$] $-2.431$ [-12.540; $3.656$]
$\omega_{33}$ $0.733$ [0.336; $1.416$] $0.770$ [0.380; $1.468$] $0.736$ [0.367; $1.339$] $0.745$ [0.343; $1.349$]
$\omega_{41}$ $18.636$ [-3.281; $44.942$] $18.428$ [-8.155; $53.426$] $14.099$ [-12.459; $42.389$] $14.149$ [-11.200; $42.412$]
$\omega_{42}$ $-37.728$ [-730.643; $916.398$] $-159.418$ [-857.598; $280.474$] $-204.434$ [-837.495; $225.430$] $-197.066$ [-804.421; $181.675$]
$\omega_{43}$ $-18.554$ [-39.966; $-0.101$] $-25.136$ [-51.927; $-4.386$] $-21.977$ [-46.403; $-3.833$] $-22.109$ [-44.681; $-2.863$]
$\omega_{44}$ $2311.940$ [1084.790; $3975.270$] $3201.290$ [1538.090; $5923.810$] $2989.335$ [1392.050; $5168.770$] $2948.195$ [1547.980; $5129.460$]
$0.95$ $\beta_{1}$ $12.566$ [12.157; $13.044$] $12.701$ [12.300; $13.108$] $12.638$ [12.221; $13.021$] $12.663$ [12.259; $13.090$]
$\beta_{2}$ $0.362$ [0.315; $0.407$] $0.338$ [0.292; $0.390$] $0.336$ [0.291; $0.383$] $0.337$ [0.295; $0.383$]
$\beta_{3}$ $7.333$ [6.626; $7.963$] $7.494$ [6.803; $8.185$] $7.325$ [6.656; $7.951$] $7.456$ [6.740; $8.075$]
$\beta_{4}$ $-0.016$ [-0.032; $0.004$] $-0.013$ [-0.030; $0.004$] $-0.012$ [-0.030; $0.006$] $-0.011$ [-0.029; $0.008$]
$\beta_{5}$ $0.006$ [-0.001; $0.013$] $0.005$ [-0.002; $0.011$] $0.005$ [-0.002; $0.011$] $0.005$ [-0.002; $0.011$]
$\sigma$ $0.020$ [0.018; $0.023$] $0.070$ [0.059; $0.080$] $0.060$ [0.042; $0.075$] $0.062$ [0.043; $0.077$]
$\gamma$ $-3.734$ [-4.940; $-2.627$] $-4.568$ [-6.571; $-2.968$] $-4.436$ [-6.410; $-2.871$]
$\alpha$ $0.012$ [0.000; $0.028$] $0.009$ [0.000; $0.026$]
$\omega_{11}$ $0.839$ [0.441; $1.318$] $1.096$ [0.501; $2.071$] $1.044$ [0.519; $1.945$] $1.059$ [0.474; $1.931$]
$\omega_{21}$ $-0.404$ [-4.923; $4.207$] $1.508$ [-6.642; $15.837$] $1.484$ [-4.387; $9.463$] $1.624$ [-8.714; $19.404$]
$\omega_{22}$ $94.944$ [28.917; $246.147$] $132.469$ [25.499; $888.207$] $128.435$ [37.715; $312.930$] $145.248$ [28.271; $1756.910$]
$\omega_{31}$ $-0.289$ [-0.623; $-0.025$] $-0.404$ [-0.941; $-0.013$] $-0.389$ [-0.913; $-0.037$] $-0.395$ [-0.937; $-0.014$]
$\omega_{32}$ $-0.165$ [-4.611; $5.048$] $-1.314$ [-11.223; $11.052$] $-1.491$ [-8.969; $3.656$] $-1.134$ [-11.285; $17.022$]
$\omega_{33}$ $0.548$ [0.268; $0.935$] $0.676$ [0.283; $1.267$] $0.658$ [0.337; $1.181$] $0.677$ [0.315; $1.305$]
$\omega_{41}$ $12.566$ [-0.376; $28.142$] $16.897$ [-4.247; $43.502$] $16.197$ [-3.971; $43.321$] $15.422$ [-7.023; $41.700$]
$\omega_{42}$ $-50.011$ [-294.766; $170.620$] $-17.611$ [-524.265; $456.264$] $-78.953$ [-497.389; $238.236$] $-89.736$ [-844.415; $747.349$]
$\omega_{43}$ $-13.750$ [-26.763; $-1.535$] $-18.478$ [-38.568; $-2.496$] $-18.885$ [-39.459; $-2.917$] $-18.189$ [-39.661; $-0.725$]
$\omega_{44}$ $1910.535$ [996.911; $3076.930$] $2276.790$ [1068.020; $3898.490$] $2522.015$ [1207.850; $4309.600$] $2596.120$ [1208.850; $4551.370$]
[…]
WEB TABLE 1: *
{ThreePartTable}
{TableNotes}
[para, flushleft]
Quantile level being modelled.
WEB TABLE 2: Posterior summaries (cGAL model, $\tau_{0}$ sensitivity). For each quantile level and each scale-inflation constant $\tau_{0}\in\left\{6.\dot{6}\dot{7},20,40\right\}$, the table lists the posterior estimate (Est.) and 95% HPD interval for all parameters. Results for higher quantiles continue on subsequent pages.
$\bm{\tau_{0}=6.\dot{6}\dot{7}}$ $\bm{\tau_{0}=20}$ $\bm{\tau_{0}=40}$
Quantilea Par. Est. 95% HPD Est. 95% HPD Est. 95% HPD
$0.05$ $\beta_{1}$ $10.765$ [10.351; $11.164$] $10.742$ [10.352; $11.186$] $10.752$ [10.325; $11.165$]
$\beta_{2}$ $0.318$ [0.279; $0.361$] $0.319$ [0.279; $0.361$] $0.318$ [0.276; $0.359$]
$\beta_{3}$ $5.600$ [4.915; $6.272$] $5.521$ [4.829; $6.234$] $5.492$ [4.820; $6.106$]
$\beta_{4}$ $-0.000$ [-0.015; $0.016$] $-0.003$ [-0.019; $0.013$] $-0.003$ [-0.019; $0.014$]
$\beta_{5}$ $0.003$ [-0.002; $0.008$] $0.004$ [-0.002; $0.010$] $0.003$ [-0.002; $0.009$]
$\sigma$ $0.055$ [0.031; $0.068$] $0.060$ [0.042; $0.073$] $0.062$ [0.049; $0.075$]
$\gamma$ $4.451$ [2.584; $7.615$] $3.590$ [2.064; $6.345$] $3.311$ [1.840; $5.420$]
$\alpha$ $0.027$ [0.005; $0.059$] $0.011$ [0.000; $0.025$] $0.007$ [0.000; $0.018$]
$\omega_{11}$ $1.110$ [0.491; $1.976$] $1.082$ [0.487; $1.991$] $1.085$ [0.477; $2.071$]
$\omega_{21}$ $2.549$ [-4.491; $14.749$] $2.982$ [-3.357; $14.280$] $3.153$ [-3.258; $13.810$]
$\omega_{22}$ $175.895$ [48.502; $527.437$] $172.263$ [54.717; $433.597$] $170.063$ [56.571; $417.020$]
$\omega_{31}$ $-0.469$ [-1.016; $-0.087$] $-0.469$ [-1.053; $-0.079$] $-0.468$ [-1.034; $-0.062$]
$\omega_{32}$ $-1.539$ [-10.618; $5.973$] $-2.292$ [-11.448; $3.306$] $-2.238$ [-10.351; $3.309$]
$\omega_{33}$ $0.728$ [0.366; $1.248$] $0.725$ [0.365; $1.285$] $0.726$ [0.309; $1.230$]
$\omega_{41}$ $12.304$ [-14.146; $40.045$] $13.343$ [-13.753; $39.236$] $14.246$ [-11.271; $43.967$]
$\omega_{42}$ $-172.218$ [-775.830; $353.575$] $-186.886$ [-761.970; $193.423$] $-169.062$ [-746.004; $174.997$]
$\omega_{43}$ $-23.256$ [-46.723; $-4.300$] $-22.979$ [-47.452; $-4.298$] $-23.993$ [-47.123; $-4.752$]
$\omega_{44}$ $3320.605$ [1533.530; $5997.970$] $3080.835$ [1480.750; $5301.480$] $3117.705$ [1525.680; $5563.090$]
$0.25$ $\beta_{1}$ $11.252$ [10.858; $11.624$] $11.239$ [10.853; $11.626$] $11.246$ [10.854; $11.618$]
$\beta_{2}$ $0.319$ [0.282; $0.360$] $0.316$ [0.281; $0.355$] $0.317$ [0.277; $0.357$]
$\beta_{3}$ $6.127$ [5.447; $6.841$] $6.110$ [5.467; $6.798$] $6.013$ [5.420; $6.685$]
$\beta_{4}$ $-0.003$ [-0.018; $0.013$] $-0.003$ [-0.020; $0.011$] $-0.005$ [-0.021; $0.010$]
$\beta_{5}$ $0.004$ [-0.002; $0.010$] $0.004$ [-0.002; $0.009$] $0.004$ [-0.001; $0.010$]
$\sigma$ $0.093$ [0.040; $0.118$] $0.102$ [0.081; $0.122$] $0.104$ [0.087; $0.122$]
$\gamma$ $0.820$ [0.220; $2.185$] $0.621$ [0.083; $1.173$] $0.564$ [0.152; $0.992$]
$\alpha$ $0.033$ [0.003; $0.077$] $0.012$ [0.001; $0.032$] $0.007$ [0.000; $0.019$]
$\omega_{11}$ $1.123$ [0.551; $2.033$] $1.119$ [0.549; $2.000$] $1.114$ [0.518; $2.051$]
$\omega_{21}$ $2.613$ [-8.442; $18.120$] $2.822$ [-12.302; $24.680$] $3.033$ [-4.894; $15.046$]
$\omega_{22}$ $217.483$ [48.891; $1076.700$] $268.924$ [46.422; $2172.440$] $205.170$ [61.611; $585.403$]
$\omega_{31}$ $-0.468$ [-1.064; $-0.085$] $-0.458$ [-1.106; $-0.031$] $-0.462$ [-1.057; $-0.087$]
$\omega_{32}$ $-1.166$ [-14.728; $13.365$] $-0.423$ [-14.681; $43.440$] $-1.910$ [-12.061; $5.103$]
$\omega_{33}$ $0.759$ [0.365; $1.380$] $0.803$ [0.325; $1.831$] $0.741$ [0.358; $1.271$]
$\omega_{41}$ $13.592$ [-14.227; $45.509$] $14.227$ [-12.733; $44.033$] $15.023$ [-10.319; $44.864$]
$\omega_{42}$ $-224.638$ [-1284.700; $481.505$] $-190.013$ [-1281.630; $1275.230$] $-222.940$ [-981.507; $207.191$]
$\omega_{43}$ $-24.804$ [-51.145; $-3.490$] $-23.840$ [-50.514; $3.075$] $-24.753$ [-49.743; $-5.291$]
$\omega_{44}$ $3377.915$ [1463.660; $6259.400$] $3254.420$ [1578.080; $6085.090$] $3186.470$ [1465.790; $5779.780$]
$0.5$ $\beta_{1}$ $11.579$ [11.190; $11.982$] $11.595$ [11.232; $11.990$] $11.597$ [11.201; $11.966$]
$\beta_{2}$ $0.318$ [0.283; $0.360$] $0.318$ [0.279; $0.358$] $0.321$ [0.282; $0.361$]
$\beta_{3}$ $6.392$ [5.773; $7.025$] $6.365$ [5.746; $7.027$] $6.381$ [5.762; $7.009$]
$\beta_{4}$ $-0.004$ [-0.019; $0.010$] $-0.006$ [-0.021; $0.011$] $-0.005$ [-0.021; $0.011$]
$\beta_{5}$ $0.004$ [-0.001; $0.010$] $0.004$ [-0.002; $0.010$] $0.004$ [-0.002; $0.010$]
$\sigma$ $0.099$ [0.029; $0.126$] $0.105$ [0.040; $0.135$] $0.108$ [0.054; $0.136$]
$\gamma$ $-0.270$ [-0.909; $0.058$] $-0.235$ [-0.870; $0.082$] $-0.229$ [-0.776; $0.112$]
$\alpha$ $0.028$ [0.002; $0.067$] $0.010$ [0.000; $0.027$] $0.007$ [0.000; $0.020$]
$\omega_{11}$ $1.115$ [0.532; $2.033$] $1.112$ [0.524; $2.038$] $1.126$ [0.545; $2.088$]
[…]